OpenStudy (anonymous):

A cookie factory uses 1/10 of a bag of flour in each batch of cookies. The factory used 1/2 of a bag of flour yesterday. How many batches of cookies did the factory make? Simplify your answer and write it as a proper fraction or as a whole or mixed number. can you please show your work too? Thanks!

OpenStudy (noseboy908):

first make 1/10 & 1/2 have the same denominator. (10)

hero (hero):

\[ \\\frac{x}{10} = \frac{1}{2} \\\frac{x}{10} = \frac{5}{10} \\x = 5 \]

OpenStudy (noseboy908):

indeed^^

OpenStudy (noseboy908):

so you use 1/10th of a bag for each batch, and you used 5/10th of a bag today, how many bags did you use?

OpenStudy (anonymous):

so they made 5 batches?

OpenStudy (anonymous):

thanks!!!! : )

OpenStudy (noseboy908):

correct, and you're welcome.